OpenStudy (thatlilmathgirl12xd):

a. Find the center of the circle. b. Find the radius. If your answer is not integer, express it in radical form. c. Write an equation for the circle.

OpenStudy (thatlilmathgirl12xd):

I apologize here is the whole thing. A diameter of a circle has endpoints P(-10,-2) and Q(4,6). a. Find the center of the circle. b. Find the radius. If your answer is not integer, express it in radical form. c. Write an equation for the circle.

OpenStudy (sshayer):

mid point is center.

OpenStudy (thatlilmathgirl12xd):

I am not very good at Geometry so I struggle with this, and my book isn't very much help when I go to look in it i end up with more questions than answers.

OpenStudy (sshayer):

find the mid point. \[x=\frac{ x _{1} +x _{2}}{ 2 }\] \[y=\frac{ y _{1}+y _{2} }{ 2 }\]

OpenStudy (thatlilmathgirl12xd):

That is the formula for midpoint?

OpenStudy (sshayer):

yes

OpenStudy (thatlilmathgirl12xd):

O.K. so then it would look like this

OpenStudy (thatlilmathgirl12xd):

-10 + 4 over 2

OpenStudy (thatlilmathgirl12xd):

and -2 + 4 over 2

OpenStudy (thatlilmathgirl12xd):

Right?

OpenStudy (sshayer):

\[y=\frac{ -2+6 }{ 2 }=?\]

OpenStudy (sshayer):

for x ,it is correct.

OpenStudy (thatlilmathgirl12xd):

oops that is what I meant.

OpenStudy (sshayer):

now what is the center.

OpenStudy (thatlilmathgirl12xd):

so the center is (-3,2)

OpenStudy (sshayer):

correct find the distance between the center and one point of diameter.

OpenStudy (sshayer):

do you know the formula for distance?

OpenStudy (thatlilmathgirl12xd):

So I use the distance formula and input (-3,2) and either (-10,-2) or (4,6)

OpenStudy (thatlilmathgirl12xd):

Correct?

OpenStudy (sshayer):

yes

OpenStudy (thatlilmathgirl12xd):

O.K. and just to make sure I remember the distance formula is

OpenStudy (thatlilmathgirl12xd):

\[\sqrt{( x _{1}- x _{2})^{2} + (y _{1}-y _{2})^{2}}\]

OpenStudy (thatlilmathgirl12xd):

Right?

OpenStudy (phi):

yes

OpenStudy (sshayer):

correct.

OpenStudy (thatlilmathgirl12xd):

So the answer would be 8.06?

OpenStudy (thatlilmathgirl12xd):

Is 8.06 my radius?

OpenStudy (sshayer):

keep it in square root form

OpenStudy (phi):

yes, but sqr(65) is more accurate

OpenStudy (thatlilmathgirl12xd):

\[\sqrt{65}\]

OpenStudy (thatlilmathgirl12xd):

Oh O.K I see

OpenStudy (sshayer):

now write the eq. of circle

OpenStudy (thatlilmathgirl12xd):

So then with making the equation I need a little extra guidance.

OpenStudy (phi):

the general form for the equation of a circle is \[ (x-h)^2 + (y-k)^2 = r^2\] where (h,k) is the center and r is the radius

OpenStudy (sshayer):

\[\left( x-h \right)^2+\left( y-k \right)^2=r^2\]

OpenStudy (thatlilmathgirl12xd):

So mine would be (x-(-3))^2 + (y-2)^2= √65 ^2

OpenStudy (sshayer):

correct. simplify it

OpenStudy (thatlilmathgirl12xd):

Simplify what the exponents?

OpenStudy (phi):

the square of a square root undoes the square root.

OpenStudy (thatlilmathgirl12xd):

Right I knew that.

OpenStudy (phi):

also, most people would simplify x- -3 to x+3

OpenStudy (thatlilmathgirl12xd):

Oh!

OpenStudy (thatlilmathgirl12xd):

Thank you very much to you both I wish I could give two Best Responses.

OpenStudy (thatlilmathgirl12xd):

:^)

OpenStudy (sshayer):

you can also write the eq. of circle from the extremities of diameter by the formula \[\left( x-x _{1} \right)\left( x-x _{2} \right)+\left( y-y _{1} \right)\left( y-y _{2} \right)=0\]

OpenStudy (thatlilmathgirl12xd):

y does it equal 0?

OpenStudy (sshayer):

yes

OpenStudy (thatlilmathgirl12xd):

Oh alright I will write that down. Thanks again!! X^D

OpenStudy (sshayer):

eq. of circle in general form is \[x^2+y^2+2gx+2fy+c=0\] where center is (-g,-f) and \[radius=\sqrt{g^2+f^2-c}\]